OpenStudy (anonymous):

Solve P = one-sixth (r + 5t) for r. You must show all work

OpenStudy (anonymous):

P=(R+5T)/6 6P=R+5T R=6P-5T

OpenStudy (anonymous):

6P=r+5t 6P-5t=r